OpenStudy (anonymous):

Find an equation for the nth term of the arithmetic sequence. a14 = -33, a15 = 9 a. an = -579 + 42(n + 1) b. an = -579 + 42(n - 1) c. an = -579 - 42(n + 1) d. an = -579 - 42(n - 1)

OpenStudy (atlas):

The general term for any arithmetic sequence is an = a + (n-1)d a14 and a15 are given - so you can find the values of a and d using them in the above equation

OpenStudy (atlas):

you can also straight away find d d = difference of any two consecutive terms of an AP= a15 - a14

OpenStudy (atlas):

so d = 9-(-33) = 42

OpenStudy (anonymous):

would it be option b? that is what i got after using the formula you gave

OpenStudy (atlas):

yeah b it is

OpenStudy (anonymous):

okay! thank you so much:)

OpenStudy (atlas):

you are welcome